If 2.131313...... is written as p/q, such that p and q are positive integers without any common factor, find value of p+q.
Answer by stanbon(75887)

If 2.131313...... is written as p/q, such that p and q are positive integers without any common factor, find value of p+q.
--------
100x = 213.131313...
x = 2.131313...
------
Subtract to get:
99x = 211
x = 211/99
--------
p/q = 211/99
p+q = 310
--------------
Cheers,
Stan H.
-------------